OpenStudy (anonymous):

-10+(-11) write in lowest terms???

OpenStudy (saifoo.khan):

-10 - 11 Same signs, add terms.. with the sign of the bigger one.. so: -21

OpenStudy (anonymous):

what about -17+4?

OpenStudy (anonymous):

how do i know if im suppose to add or subract? its confusing!

OpenStudy (saifoo.khan):

different signs, subtract them.. with the sign of bigger one: -13

OpenStudy (saifoo.khan):

the signs tell you. If you have same signs: Add them. Different signs: subtract them. In both cases, put the sign of the bigger term.

OpenStudy (anonymous):

5-(-3)? is this gonna be 8?

OpenStudy (anonymous):

im so confused :(

OpenStudy (saifoo.khan):

great, right.

OpenStudy (anonymous):

im right???!!!

OpenStudy (saifoo.khan):

You get a medal for that answer!! :D yes, u r correct.

OpenStudy (anonymous):

yay!!! :D okay this one -32 - (-12) = 44????

OpenStudy (saifoo.khan):

Damn it, no! :/

OpenStudy (saifoo.khan):

solve step by step.

OpenStudy (anonymous):

aw :( but you said if there the same signs to add them?

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

YUP

OpenStudy (saifoo.khan):

Right. but look: -32 + 12 different signs, subtract them. 20. BUT WITH THE SIGN OF THE BIGGER NUMBER. so, -20

OpenStudy (saifoo.khan):

kahn, sorry u r wrong.

OpenStudy (anonymous):

so if its (-12) im suppose to make it the opposite so it will turn to a positive?

OpenStudy (saifoo.khan):

No, - * - = +

OpenStudy (anonymous):

-32-(-12) 32+(-12) 32-12=20 or i guess wt u r saying it would be like , -32-(-12) -32+12=-20 (sign of bigger value which is 32)

OpenStudy (saifoo.khan):